# GETReason: Enhancing Image Context Extraction through Hierarchical Multi-Agent Reasoning

### v2

@@ Abstract
Shikhhar Siingh Abhinav Rawat Chitta Baral  Vivek Gupta Arizona State University {ssiingh, arawat11, chitta, vgupt140}@asu.edu
**footnotetext: These authors contributed equally to this work and share first authorship.footnotetext: This author supervised the research and serves as the corresponding author.
@@ 1 Introduction
Public event images are rich sources of contextual information. They capture moments of societal importance, reflecting not just visual content but also the broader narratives they represent. Extracting such context enhances news reporting, sentiment analysis, and archival work.
Despite their relevance, event-centric image analysis remains underexplored. Existing methods often fall short in capturing nuanced event-specific information, as shown in Table 2. Standard image captioning models, particularly encoder-decoder approaches, focus on surface features—objects, people, and actions—while missing deeper context. Even advanced models Li et al. (2022); Hu et al. (2023) mainly describe what is seen, without identifying the event’s geopolitical or historical significance.
@@
Visual Language Models (VLMs) Lu et al. (2019) offer strong potential for this task, yet current methods often fail to extract the rich, pre-trained knowledge within these models. For instance, Figure 1 shows President Joko Widodo’s second-term inauguration on October 20, 2019, a politically charged moment. A typical captioning model might note the speaker, the reverend, and background figures, but miss the event’s significance—understandable only by synthesizing these details.
Reasoning-based methods Zhang et al. (2024); Song et al. (2024) can infer generic scenarios like ceremonies but rarely capture the full historical or political context. Retrieval-Augmented Generation (RAG) Lewis et al. (2020) integrates external knowledge for better context but is prone to misinformation and noisy sources Zeng et al. (2025); Deng et al. (2024), limiting its reliability for precise event interpretation. Recent advances in multi-agent frameworks Dinh and Chan (2025); Ghafarollahi and Buehler (2024); Wang et al. (2025); Ng et al. (2024) show that distributing reasoning tasks across specialized agents improves contextual understanding across domains. Inspired by this divide-and-conquer paradigm, GETReason applies multi-agent reasoning to a large, diverse event image dataset—enhancing extraction accuracy, robustness, and generalizability.
Motivation: The challenges above underscore why extracting event-specific context from public event images remains difficult. Existing approaches either limit themselves to shallow visual cues or depend on external retrieval, which risks hallucinations, misinformation, and fragmented context. True event understanding requires reasoning about temporal, spatial, and sociopolitical factors—often implicit rather than visually explicit.
Addressing this problem is vital for domains like journalism, archival research, sentiment analysis, and historical documentation, where context matters beyond visual recognition. A robust solution would enable the automatic construction of richer, more accurate narratives around public events, enhancing both human interpretation and downstream machine applications.
Main Contributions: Our approach is driven by the need for a systematic, reasoning-based method to extract three key components from an image: (1) geospatial location (e.g., "Indonesia"), (2) temporal information (e.g., "October 20, 2019"), and (3) event-specific details (e.g., "inauguration of President Joko Widodo’s second term amid public tensions"). To overcome prior limitations, we introduce a structured, multi-agentic framework where specialized agents collaboratively extract and validate contextual elements. Inspired by Ng et al. (2024), this architecture reduces errors and hallucinations through cross-agent validation, improving both precision and reliability. A thorough working example for the framework is explained in Appendix B under Working Example.
In summary, our key contributions are as follows:
@@ 2.1 TARA*
TARA Event Information Augmentation.
@@
2.2 WikiTiLo*
@@
Deduction Augmentation.
Finally, we used the augmented ground truths as references for the images on the VLM to generate the deduction inferences and reasoning from the images for all the informations present in the augmented dataset. For example, in Figure 1, the event deduction of the image will be: The image shows Joko Widodo taking an oath on the Quran during his initiation ceremony, the temporal deduction of the image will be: the image is about the initiation ceremony of Joko Widodo which took place in 20 October, 2019 and the geospatial ceremony of the image will be: from the attire of the people in the background and the setting, it looks like a government setting in Jakarta. Since the event is inauguration of Joko Widodo, the geospatial location is at the People’s Consultative Assembly (MPR) Building.
@@ 3.3 Multi-Agentic Extraction
More advanced metrics such as ROUGE  Lin (2004), BLEU  Papineni et al. (2002), and METEOR  Banerjee and Lavie (2005) improve evaluation but remain biased toward word overlap and order, limiting their ability to capture meaning. ROUGE also favors longer texts  Lin (2004), while CIDEr  Vedantam et al. (2015), though effective for image captioning, is not designed for reasoning tasks requiring spatio-temporal evaluation. Human evaluation, while useful for qualitative assessment, lacks consistency for quantitative analysis.
For geospatial and temporal evaluation, a simple matching based F1 scoring can be used like in Fu et al. (2022), however, a major drawback of this metric is that it does not score based on the reasoning ability of the model, i.e. the closeness of the value to the ground truth is not measured. Thus, to address all these problems, we designed a new metric - GREAT. This metric comprises of three parts -
@@
Let $E_{i}$ and $B_{i}$ denote the augmented event and background for image $i$ and let $e_{i}$ and $b_{i}$ denote the corresponding predicted event and background values. The concatenated event and background representations are given by $e_{i}+b_{i}$ for predictions and $E_{i}+B_{i}$ for ground truth labels.
We first compute the cosine similarity between the embeddings generated by a Sentence Transformer model from HuggingFace Wolf et al. (2019), specifically the all-mpnet-base-v2 model, for the predicted and ground truth concatenated event-background pairs. Since the cosine similarity lies within the range $[-1,1]$, we shift this range to $[0,1]$ by using the formula:
@@
By leveraging cosine similarity, GREAT measures the semantic similarity between the predicted and ground truth event-background pairs, effectively reflecting the model’s reasoning ability.
@@
To assess the accuracy of predicted geospatial locations, we define a similarity metric that evaluates the hierarchical correctness of location predictions and computes a distance-based score using the Haversine formula. Given a ground truth location $gt$ and a predicted location $pred$, we first determine the deepest available location level among country, state/province, and city. If no valid prediction exists, the score is set to 0.
The geospatial similarity is then computed based on the great-circle (Haversine) distance Hamilton (19th Century) between the latitude-longitude coordinates of the ground truth and predicted locations. The Haversine distance $d$ is given by:
@@ 5.1 Experimental Settings
For GPT-4o-minihttps://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/, we utilized the batch API. While OpenAI’s models do not directly reveal public figureshttps://openai.com/policies/usage-policies/, they can ground them if their identities are provided externally in the prompt. To leverage this, we used the CelebrityFaces API from AWS Rekognitionhttps://aws.amazon.com/rekognition/ to identify public figures in the dataset images with a confidence score of at least 99%. These identified public figures were then explicitly included in the prompts during the experiments to enable more precise entity recognition.
For Qwen-VL2.5-7B-Instruct Team (2025b)  Wang et al. (2024)  Bai et al. (2023), the experiments were conducted on three NVIDIA A100 Tensor Core GPUs, each with 80 GB of GPU memory, alongside 8-core CPUs with 32 GB of RAM. These resources were hosted on the Arizona State University’s Sol supercomputer Jennewein et al. (2023)
@@
To evaluate the reasoning capabilities of our framework, we compared the reasoning output generated by our model to the ground truth across three modalities on the TARA dataset on Gemini 1.5 pro-002, utilizing the Event Evaluation metric from GREAT.
5.2 Main Results.
Table 2 presents the performance comparison of our model’s three variants against 11 baseline strategies including a zeroshot Chain-of-Thoughts prompt and a Detective Agent prompt across three information types. For RAG based approaches, we modified the RAG parts of the prompts to using the pretrained knowledge base instead of internet extraction. Evaluating on the GREAT evaluation metric, we observe that across almost all cases, our framework consistently achieves the best performance for all reasoning types viz. event, temporal, and geospatial, demonstrating its superiority over other baselines. Notably, Good Guesser also performs well in geospatial reasoning, surpassing even our framework 3 out of 6 times and achieving the second highest scores in most of the other tasks.
These results highlight the robustness of GETReason while also indicating that the Good Guesser baseline remains a strong contender in geospatial reasoning. A decline in our framework’s performance is observed when evaluated with GPT-4o-mini. This can be attributed to the comparatively lower reasoning capability of GPT-4o-mini in handling complex tasks compared to the other two models as well as its more advanced counterpart GPT-4o.
Analysis. We observe that GETReason achieves scores of 81.4%, 76.9% and 70.2% for geospatial, temporal and event deduction respectively. The consistently high scores achieved by individual reasoning instances suggest a strong likelihood that the model accurately extracts information from images through effective reasoning.
5.3 Ablation Analysis.
@@
5.4 Multi-Dimensional Error Analysis
@@
Table 4: Net percentage gain in samples where GETReason outperforms individual baselines for TARA with Gemini 1.5 Pro-002. GTS = Generate then Select. DA = Detective Agent
@@
5.5 Across model analysis
Along the results obtained from the 3 models, Gemini 1.5 pro-002, GPT-4o-mini and Qwen2.5VL-7B, we found that while all three models perform competitively at extracting the event reasoning with Gemini leading among them, Gemini performs better at geospatial and temporal extraction tasks. Following Gemini, among Qwen and GPT, Qwen consistently performs better at geospatial extraction, while GPT performs better at temporal extraction.
Analysis Gemini, being much bigger than the other two models performs much better at reasoning, while Qwen and GPT-4o-mini being much smaller than Gemini and almost similar in size to each other, perform similarly outperforming each other in alternate tasks.,
5.6 Effectiveness Analysis
There are several compelling reasons why our methodology is well-suited to this problem and consistently outperforms both existing baselines and those developed by us. These reasons can be broadly categorized into two groups:
1. Fundamental methodological differences:
a. Hierarchical Multi-Agent Framework:
Our method employs a hierarchical multi-agent framework, reducing cognitive load on the VLM, minimizing hallucinations, and enabling context resetting—unlike baselines that rely on single, monolithic prompts.
b. Structured Prompting:
We use structured, role-specific prompts and controlled information flow, providing greater consistency and interpretability than baselines that depend on ad-hoc prompt engineering.
c. Schema-Driven Reasoning:
By combining structured schema extraction with flexible reasoning, our approach achieves both precision and adaptability, addressing the inefficiencies and hallucinations often seen in less structured methods.
d. Systematic Cue Integration:
Our framework explicitly handles temporal, geospatial, and event-specific cues, which are often overlooked or inconsistently extracted in prior methods (see Section 6).
e. Robust, Large-Scale Evaluation:
GETReason is evaluated on over 17,000 examples, enabling a more robust and generalizable assessment than prior works limited to datasets of around 2,000 samples.
2. Diverse Baseline Limitations:
As detailed in Section 6, baseline methods span object detection, shallow reasoning, and retrieval-based approaches—each introducing specific limitations in contextual depth, prompt precision, or overreliance on external sources.
@@ 6 Comparison with Related Work
Given the diversity of these baselines, we ensured a fair comparison by aligning all methods to our evaluation schema and restricting them to the VLM’s internal knowledge base.
@@ Acknowledgements
We thank the anonymous reviewers for their helpful feedback. We also thank Zenia Shaikh for her valuable assistance and contributions to this work. We thank the Complex Data Analysis and Reasoning Lab at Arizona State University for computational support, and Krishna Singh Rajput for his support during the submission process.
@@ C.1 Additional Ablation Results
Further Discussion
In this ection we discussion error propogation in GETReason, RAG vs GETReason, and usability and scalability of GETReason.
C.2 Error propagation
@@ C.3 RAG vs GETReason
Additionally, we sincerely apologize for the missing citations and typos. We appreciate your review, and we will ensure that the final version is free from any form of errors, confusing sections, or ill-supported claims. We have identified areas of improvement and have started working on them.
Finally, we truly appreciate this opportunity to clarify these aspects, and we will ensure that they are more explicitly stated in the final version of the paper.
C.4 Usability and Scalability of GETReason
@@
By leveraging these two levels of parallelization, our approach effectively balances processing speed and output quality, ensuring scalability without compromising accuracy. We will add all these details in the appendix.

### v3

@@ Abstract
Shikhhar Siingh Abhinav Rawat Chitta Baral  Vivek Gupta School of Computing and Augmented Intelligence, Arizona State University {ssiingh, arawat21, chitta, vgupt140}@asu.edu
**footnotetext: These authors contributed equally to this work.footnotetext: This corresponding author supervised the research.
@@ 1 Introduction
Public event images capture moments of societal and historical importance—presidential inaugurations, mass protests, international summits—serving not only as visual records but as rich contextual artifacts. Understanding such images requires more than surface-level description; it demands inference about geopolitical, temporal, and event-specific factors often implicit in the scene. Despite their significance for journalism, archival analysis, and public discourse, event-centric image understanding remains underexplored.
@@
Most existing vision-language methods fall short in this setting. Traditional captioning models and encoder-decoder architectures describe visible content—objects, people, and actions—but fail to infer deeper meaning. Even advanced visual-language models Li et al. (2022); Hu et al. (2023) tend to narrate what is seen, missing why it matters. For instance, given an image of President Joko Widodo’s inauguration (Figure 1), these models might mention key figures but omit the event’s political gravity and temporal context. However, these methods fall short of true event understanding, as they fail to reason about temporal, spatial, and sociopolitical factors—many of which are implicit rather than directly observable.
Efforts to move beyond superficial cues include reasoning-based methods Zhang et al. (2024); Song et al. (2024), which can infer coarse event types like “ceremony,” and retrieval-augmented generation (RAG)Lewis et al. (2020), which brings in external knowledge. Yet both have limitations: the former lacks specificity, and the latter is vulnerable to hallucinations and misinformation due to noisy sourcesZeng et al. (2025); Deng et al. (2024). These approaches often miss the nuanced, high-precision understanding required for real-world applications. Therefore, a robust solution is needed to automatically generate richer and more accurate narratives of public events, enhancing both human comprehension and downstream machine applications.
Recent advances in multi-agent frameworks Dinh and Chan (2025); Ghafarollahi and Buehler (2024); Wang et al. (2025); Ng et al. (2024) demonstrate that dividing reasoning tasks among specialized agents can significantly improve contextual understanding across domains. However, these methods have primarily focused on textual or code-based settings. This raises an important question: Can a structured multi-agent framework effectively support contextual reasoning over complex visual data, such as public event images?
To answer this question, we introduce GETReason, a structured multi-agent framework for extracting rich, contextual narratives from public event images. GETReason decomposes the task into three core sub-problems, each handled by a dedicated agent: (1) geospatial inference (e.g., “Indonesia”), (2) temporal inference (e.g., “October 20, 2019”), and (3) event-specific interpretation (e.g., “inauguration of President Joko Widodo’s second term amid political unrest”). Building on the divide-and-conquer principles of recent multi-agent systems Ng et al. (2024); Dinh and Chan (2025), GETReason enables collaborative cross-agent validation to enhance factual accuracy, reduce hallucinations, and improve generalization across diverse event types.
We evaluate GETReason on a large-scale dataset of public event images and demonstrate substantial gains over existing captioning and reasoning baselines. Our results highlight the importance of structured, role-specialized reasoning in achieving precise and reliable event understanding—unlocking new capabilities for downstream applications in news generation, historical archiving, and multimodal analysis. In summary, our key contributions are as follows:
@@
2 Task Description
We define the task of event-centric image understanding as generating a structured triplet from a single image depicting a public event. Given an image $I$, the model must infer: (1) a location $L$ (e.g., city or country), (2) a temporal reference $T$ (e.g., specific date or period), and (3) an event description $E$ that captures the scene’s sociopolitical significance. Unlike standard captioning, which describes visible entities or actions, this task requires reasoning about partially implicit geopolitical, temporal, and narrative cues.
In Figure 1, for instance, the model links the image of President Joko Widodo to an oath-taking ceremony, the green-bound 1945 Constitution, and the presence of key political figures to derive the structured output: Expected Output: $L$: “Jakarta, Indonesia”; $T$: “October 20, 2019”; $E$: “Inauguration of President Joko Widodo’s second term amid protests.”
Each component involves distinct reasoning challenges: $L$ depends on geographic signals (e.g., flags, landmarks); $T$ draws on visible dates, seasonal features, or known timelines; and $E$ requires synthesizing the visual context with latent knowledge of political and historical events. This formulation enables structured, interpretable outputs for applications such as news analysis, archival indexing, and societal trend monitoring, while posing challenges in multimodal inference and hallucination control.
@@ 3.1 TARA*: TARA Enhancement
Spatio-Temporal Augmentation.
Event Information Augmentation.
@@
Deduction Augmentation.
Finally, we used the augmented ground truths as references for the images on the VLM to generate the deduction inferences and reasoning from the images for all the information present in the augmented dataset. For example, in Figure 1, the event deduction of the image will be: The image shows Joko Widodo taking an oath on the Quran during his initiation ceremony, the temporal deduction of the image will be: the image is about the initiation ceremony of Joko Widodo which took place in 20 October, 2019 and the geospatial ceremony of the image will be: from the attire of the people in the background and the setting, it looks like a government setting in Jakarta. Since the event is inauguration of Joko Widodo, the geospatial location is at the People’s Consultative Assembly (MPR) Building.
3.2 WikiTiLo*: WikiTiLo Enhancement
@@ 4.3 Multi-Agentic Extraction
More advanced metrics such as ROUGE  Lin (2004), BLEU  Papineni et al. (2002), and METEOR  Banerjee and Lavie (2005) improve evaluation but remain biased toward word overlap and order, limiting their ability to capture meaning. ROUGE also favors longer texts  Lin (2004), while CIDEr  Vedantam et al. (2015), though effective for image captioning, is not designed for reasoning tasks requiring spatio-temporal evaluation. Human evaluation, while useful for qualitative assessment, lacks quantitative consistency.
For geospatial and temporal evaluation, a simple matching based F1 scoring can be used like in Fu et al. (2022), however, a major drawback of this metric is that it does not score based on the reasoning ability of the model, i.e. the closeness of the value to the ground truth is not measured. Thus, to address these issues, we propose a new metric - GREAT, which comprises of three parts:
@@
Let $E_{i}$ and $B_{i}$ denote the augmented event and background for image $i$ and let $e_{i}$ and $b_{i}$ denote the corresponding predicted event and background values. The concatenated event and background representations are given by $e_{i}+b_{i}$ for predictions and $E_{i}+B_{i}$ for ground truth labels. We first compute the cosine similarity between the embeddings generated by a Sentence Transformer model from HuggingFace Wolf et al. (2019), specifically the all-mpnet-base-v2 model, for the predicted and ground truth concatenated event-background pairs. Since the cosine similarity lies within the range $[-1,1]$, we shift this range to $[0,1]$ by using the formula:
@@
By leveraging cosine similarity, GREAT measures the semantic similarity between the predicted and ground truth event-background pairs, effectively evaluating model’s reasoning strength.
@@
To assess the accuracy of predicted geospatial locations, we define a similarity metric that evaluates the hierarchical correctness of location predictions and computes a distance-based score using the Haversine formula. Given a ground truth location $gt$ and a predicted location $pred$, we first determine the deepest available location level among country, state/province, and city. If no valid prediction exists, the score is set to 0. The geospatial similarity is then computed based on the great-circle (Haversine) distance Hamilton (19th Century) between the latitude-longitude coordinates of the ground truth and predicted locations. The Haversine distance $d$ is given by:
@@ 6 Experiments
For GPT-4o-minihttps://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/, we utilized the batch API. While OpenAI’s models do not directly reveal public figureshttps://openai.com/policies/usage-policies/, they can ground them if their identities are provided externally in the prompt. To leverage this, we used the CelebrityFaces API from AWS Rekognitionhttps://aws.amazon.com/rekognition/ to identify public figures in the dataset images with a confidence score of at least 99%. These identified public figures were then explicitly included in the prompts during the experiments to enable more precise entity recognition. For Qwen-VL2.5-7B-Instruct Team (2025b)  Wang et al. (2024)  Bai et al. (2023), the experiments were conducted on three NVIDIA A100 Tensor Core GPUs, each with 80 GB of GPU memory, alongside 8-core CPUs with 32 GB of RAM. These resources were hosted on the Arizona State University’s Sol supercomputer Jennewein et al. (2023)
@@
To evaluate the reasoning capabilities of our framework, we compared the reasoning output generated by our model to the ground truth across three modalities on the TARA dataset on Gemini 1.5 pro-002, utilizing the Event Evaluation metric from proposed GREAT metric.
6.1 Results and Analysis.
Table 2 presents the performance comparison of our model’s three variants against 11 baseline strategies including a zeroshot Chain-of-Thoughts prompt and a Detective Agent prompt across three information types. For RAG based approaches, we modified the RAG parts of the prompts to using the pretrained knowledge base instead of internet extraction. Evaluating on the GREAT evaluation metric, we observe that across almost all cases, our framework consistently achieves the best performance for all reasoning types viz. event, temporal, and geospatial, demonstrating its superiority over other baselines. Notably, Good Guesser also performs well in geospatial reasoning, surpassing even our framework 3 out of 6 times and achieving the second highest scores in most tasks.
These results highlight the robustness of GETReason while also indicating that the Good Guesser baseline remains a strong contender in geospatial reasoning.
Across Models Analysis
Across the three models—Gemini 1.5 Pro-002, GPT-4o-mini, and Qwen2.5VL-7B—all perform competitively in event reasoning, with Gemini leading overall. Gemini also outperforms the others in geospatial and temporal extraction. Between Qwen and GPT, Qwen consistently excels in geospatial tasks, while GPT performs better in temporal tasks. Given its larger size, Gemini shows stronger reasoning capabilities, while the smaller and similarly sized Qwen and GPT-4o-mini perform comparably, each outperforming the other in specific tasks.
Deduction Analysis
In order to evaluate our framework’s reasoning capabilities, we compared its output to ground truth across three modalities on the TARA dataset using Gemini 1.5 Pro-002 and the Event Evaluation metric from GREAT. GETReason achieves scores of 81.4%, 76.9%, and 70.2% for geospatial, temporal, and event deduction, respectively. These consistently high scores across modalities indicate the model effectively extracts information from images through accurate reasoning.
Ablation Analysis.
@@
Multi-Dimensional Error Analysis.
@@
Table 4: Net percentage gains where GETReason outperforms individual baselines for TARA with Gemini 1.5 Pro-002. GTS = Generate then Select. DA = Detective Agent
@@
6.2 GETReason Effectiveness
There are two primary reasons GETReason consistently outperforms both existing and internally developed baselines:
(1) Methodological Advances.
Our approach introduces several core innovations: (a) Hierarchical Multi-Agent Design: By decomposing the task across specialized agents, our method reduces VLM overload, resets context between subtasks, and mitigates hallucinations—unlike monolithic prompting schemes. (b) Structured Prompting: We employ role-specific prompts with controlled information flow, yielding more consistent and interpretable outputs compared to ad-hoc prompting strategies. (c) Schema-Guided Reasoning: Our framework blends schema extraction with flexible reasoning, enabling precision without sacrificing adaptability—addressing common limitations of unstructured methods. (d) Systematic Cue Integration: Temporal, spatial, and event-specific cues are explicitly handled, unlike prior models that overlook or inconsistently process such signals (see Section 7). (e) Scalable Evaluation: GETReason is evaluated on over 17,000 public event images, thus a more reliable assessment than prior work typically benchmarked on $\sim$2,000 samples.
(2) Baseline Limitations.
Competing methods—spanning object-centric captioning, shallow reasoning, and retrieval-augmented generation—struggle with contextual grounding, prompt coherence, and resilience to misinformation. Our structured, agent-based design addresses these issues holistically.
@@ Acknowledgements
We thank the anonymous reviewers for their helpful feedback. We also thank Zenia Shaikh for her valuable assistance and contributions to this work. We thank the Complex Data Analysis and Reasoning Lab at Arizona State University for computational support, and Krishna Singh Rajput for his support during the submission process. Lastly, thanks to our lab cat, Coco, for keeping our professor company and adding some much-needed humor during deadline crunches.
@@ C.1 Additional Ablation Results
Appendix D Further Discussion
In this section we discussion error propagation in GETReason, RAG vs GETReason, and usability and scalability of GETReason.
D.1 Error propagation
@@ D.2 RAG vs GETReason
D.3 Usability and Scalability of GETReason
@@
By leveraging these two levels of parallelization, our approach effectively balances processing speed and output quality, ensuring scalability without compromising accuracy.